# Quantifying and Alleviating Co-Adaptation in Sparse-View 3D Gaussian Splatting

### v2

@@ A.9 Exploration of Advanced Dropout Strategies
Concrete Dropout. We implement a variant of Concrete Dropout [Gal et al., 2017] by learning a per-Gaussian dropout probability $p_{i}\in(0,1)$. During training, we sample a soft mask $z_{i}$ using the Binary Concrete distribution:
@@ A.10 Exploration of Multi-View Learning
To further explore the impact of multi-view learning under sparse-view 3DGS settings, we implemented the official MVGS [Du et al., 2024] using 3 input views. As shown in Table 14, the MVGS model yields significantly degraded performance across all image quality metrics under sparse-view input, suggesting its lack of adaptation to sparse supervision.
@@ A.11 Extended Experimental Details
To ensure fair comparisons, we adopt unified parameter settings across all scenes within each dataset. For example, while the original DNGaussian [Li et al., 2024] uses different training configurations for each Blender scene, we apply a consistent setup across all scenes. Binocular3DGS is retrained using white backgrounds to match other methods, as its original results were obtained using black backgrounds. For co-adaptation suppression, we use a fixed dropout probability of 0.2 across all methods and datasets, based on ablations conducted on Binocular3DGS. Because each method learns a distinct opacity distribution, we tune the opacity noise scale individually for each method within $[0.05,0.8]$ and fix it across all scenes in the same dataset. For CA score computation, we randomly drop 50% of the contributing Gaussians for each view and calculate the pixel-wise variance across multiple renderings. The number of training iterations for each baseline follows its official implementation.

### v3

@@ A.9 Exploration of Advanced Dropout Strategies
Concrete Dropout. We implement a variant of Concrete Dropout [59] by learning a per-Gaussian dropout probability $p_{i}\in(0,1)$. During training, we sample a soft mask $z_{i}$ using the Binary Concrete distribution:
@@ A.10 Exploration of Multi-View Learning
To further explore the impact of multi-view learning under sparse-view 3DGS settings, we implemented the official MVGS [60] using 3 input views. As shown in Table 14, the MVGS model yields significantly degraded performance across all image quality metrics under sparse-view input, suggesting its lack of adaptation to sparse supervision.
@@ A.11 Extended Experimental Details
To ensure fair comparisons, we adopt unified parameter settings across all scenes within each dataset. For example, while the original DNGaussian [2] uses different training configurations for each Blender scene, we apply a consistent setup across all scenes. Binocular3DGS is retrained using white backgrounds to match other methods, as its original results were obtained using black backgrounds. For co-adaptation suppression, we use a fixed dropout probability of 0.2 across all methods and datasets, based on ablations conducted on Binocular3DGS. Because each method learns a distinct opacity distribution, we tune the opacity noise scale individually for each method within $[0.05,0.8]$ and fix it across all scenes in the same dataset. For CA score computation, we randomly drop 50% of the contributing Gaussians for each view and calculate the pixel-wise variance across multiple renderings. The number of training iterations for each baseline follows its official implementation.